# CityLens: Benchmarking Large Language-Vision Models for Urban Socioeconomic Sensing

### v1

@@ 6 Discussion
Our benchmark results highlight both the potential and limitations of using LLVMs to predict region-level socioeconomic indicators. While some tasks, particularly those with visually salient correlates such as building height achieve moderate performance, most indicators remain challenging to estimate accurately. Their outputs often converge around city-level averages, suggesting that the model lacks sensitivity to intra-city variation, and consequently exhibits limited geospatial grounding. In the two LLVM as predictor paradigms, we observe that different tasks tend to favor different strategies: some perform better under direct value prediction, while others benefit from normalized estimation. Overall, our results indicate that the Feature-Based Regression paradigm, where the LLM functions as a feature enhancer, significantly outperforms the two predictor-based methods. These findings suggest several promising directions for future research. First, while the feature-based method relies on a trained LASSO regressor, the predictor-based methods are evaluated in a zero-shot setting. This highlights the potential benefit of fine-tuning LLVMs directly for socioeconomic indicator prediction tasks. Second, improvements in prompt engineering could enhance model performance. In particular, incorporating Chain-of-Thought (CoT) prompting may help guide the model through intermediate reasoning steps, yielding more reliable numerical estimates. Finally, we envision the development of a domain-specific agent framework tailored to urban socioeconomic sensing, which could combine visual perception, geospatial knowledge, and reasoning modules to make robust and context-aware predictions in real-world scenarios.
Appendix A Technical Appendices and Supplementary Material
A.1 Summary of Indicators in CityLens
@@
Under the economy domain, we cover 7 critical indicators: Gross Domestic Product (GDP), house price, population, median household income, poverty 100%, poverty 200% and income gini coefficient. For GDP, we utilize a global dataset that provides GDP estimates with a spatial resolution of 1 km [39]. For population, we adopt estimates from WorldPop [37], a global demographic dataset with 1 km spatial resolution that provides consistent population counts across countries. For house price, we collect data from multiple sources tailored to each country’s context: (1) For US cities, we use the Zillow Home Value Index (ZHVI) [50], available at the ZIP code level. We map these values to census tract boundaries using spatial overlays, enabling fine-grained local prediction. (2) For UK cities, we obtain house price data from [16], targeting the Middle Layer Super Output Area (MSOA) level. (3) For Chinese cities, we collect house price data from LianJia [34], one of China’s largest online real estate platforms. For median household income, poverty 100%, and poverty 200%, we obtain the raw data from [1]. We obtain the ground-truth values for the income Gini coefficient from [46].
@@
In the transport domain, we include seven indicators: PMT, VMT, PTRP, VTRP, walk and bike ratio, public transport ratio, and drive ratio, following the design of [14]. The underlying data is sourced from [36], which provides commuting behavior statistics at the census tract level across the United States. Table 3 outlines the definitions of the seven transport related indicators.
@@
In the crime domain, we focus on two indicators: violent crime incidence and non-violent crime incidence, both defined as the number of crime occurrences per census tract. The data is collected from the official websites of individual US cities [12, 30, 11], which publish annual crime reports and geolocated incident-level data.
@@
For the health domain, we include 9 kinds of indicators to capture different dimensions of urban health outcomes: obesity, diabetes, cancer, no leisre-time physical activity (LPA), mental health, physical health, depression rate, accessibility to healthcare, and life expectancy. The first 7 tasks focus on the United States only, using data from “Local Data for Better Health” [9]. The accessibility to healthcare task is defined globally, using a dataset that quantifies walking-only travel time to the nearest healthcare facility [41]. The life expectancy task targets the United Kingdom, where we use data from [16] to obtain male life expectancy estimates at the MSOA level.
@@
In the environment domain, we consider two indicators: carbon emissions and building height. For carbon, we use global estimates from [31]. We use global building height data obtained from [33], which provides global coverage at a spatial resolution of 100 meters.
@@
Following [27], we use the bachelor ratio, defined as the proportion of residents holding a bachelor’s degree or higher, as the target variable in the education domain. The ground-truth data for this indicator is obtained from SafeGraph [1], which provides fine-grained demographic datasets across the United States.
@@
Economy GDP [39]
House Price [50, 16, 34]
Population [37]
Median Income [1]
Poverty 100% [1]
Poverty 200% [1]
Income Gini Coefficient [46]
Education Bachelor Ratio [1]
Crime Violent [12, 30, 11]
Non-Violent [12, 30, 11]
Transport PMT [36]
VMT [36]
PTRP [36]
VTRP [36]
Walk and Bike [36]
Drive Ratio [36]
Public Transport [36]
Health Obesity [9]
diabetes [9]
LPA [9]
Cancer [9]
Mental Health [9]
Physical Health [9]
Depression Rate [9]
Life Expectancy [16]
Accessibility to Healthcare [41]
Environment Carbon Emissions [31]
Building Height [33]
@@
A.2 Prompt Design and Case Analysis
We provide additional insights into the design and behavior of LLVMs under the three evaluation methodologies. Table 4 summarizes the representative prompts used in the three paradigms, highlighting their structural differences and role-setting strategies. Figure 6 shows a case in the Direct Metric Prediction setting, where the model refuses to estimate GDP due to insufficient information. Figure 7 shows an example under the Normalized Estimation setting, where the model is asked to predict the bachelor ratio based on regional images. Figure 8 presents an example from the Feature-Based Regression method, where the model scores a street view image along 13 predefined urban visual attributes to support downstream prediction.
@@ A.2 Prompt Design and Case Analysis
Figure 7: Prompt example for Normalized Metric Estimation.
Figure 8: Prompt template for guiding large language-vision models to extract 13 visual features from a street view image.
A.3 Supplementary Comparison of Three Evaluation Methodologies
@@
From the figures, it is evident that the Feature-Based approach, where the large language-vision model acts as a feature enhancer, consistently outperforms the Direct and Normalized approaches, in which the model is expected to behave as a numerical predictor. This suggests that current LLVMs, while powerful in perceptual and language tasks, are still more effective when used to extract structured visual representations rather than to directly generate precise socioeconomic estimates. Although large language-vision models have made impressive strides, accurately predicting fine-grained, region-level socioeconomic indicators remains highly challenging, highlighting the need for further advancements.

### v2

@@ A.2 Discussion
Our benchmark results highlight both the potential and limitations of using LVLMs to predict region-level socioeconomic indicators. While some tasks, particularly those with visually salient correlates such as building height achieve moderate performance, most indicators remain challenging to estimate accurately. Their outputs often converge around city-level averages, suggesting that the model lacks sensitivity to intra-city variation, and consequently exhibits limited geospatial grounding. In the two LVLM as predictor paradigms, we observe that different tasks tend to favor different strategies: some perform better under direct value prediction, while others benefit from normalized estimation. Overall, our results indicate that the Feature-Based Regression paradigm, where the LLM functions as a feature enhancer, significantly outperforms the two predictor-based methods. These findings suggest several promising directions for future research. First, while the feature-based method relies on a trained LASSO regressor, the predictor-based methods are initially evaluated in a zero-shot setting. Through fine-tuning LVLMs, we have demonstrated the significant potential of training domain-specific LVLMs for urban socioeconomic sensing. Second, future improvements may come from designing prompts that more closely reflect human reasoning patterns, beyond standard CoT prompting. While our current experiments already incorporate CoT prompts, we believe that further performance gains may be achieved through cognitively informed prompt design. To support this direction, we also outline a hypothesis in Appendix A.10 on how LVLMs accomplish urban socioeconomic sensing. Finally, we envision the development of a domain-specific agent framework tailored to urban socioeconomic sensing, which could combine visual perception, geospatial knowledge, and reasoning modules to make robust and context-aware predictions in real-world scenarios.
A.3 Benchmark Results with Mapillary Street View Images
To promote transparency, completeness, and reproducibility, we construct an alternative version of CityLens using publicly available street view images from the open-source Mapillary platform, referred to as CityLens-Mapillary. This version serves as an open-source complement to the main benchmark and is intended for use in scenarios where proprietary street view APIs (Google and Baidu) are inaccessible, representing a worst-case scenario due to access restrictions or licensing limitations.
Due to the relative sparsity and inconsistency of open-source street view coverage, the number of supported tasks in CityLens-Mapillary is reduced. Table 3 summarizes the number of available prediction tasks for each indicator, comparing the original CityLens with the Mapillary-based version. Table 4 presents the results on CityLens-Mapillary. We observe a modest drop in performance compared to the original CityLens, which we attribute to the generally lower image quality and coverage of the Mapillary platform. For instance, in the UK region, predictions for Life Expectancy show several negative $R^{2}$ values, likely resulting from degraded visual input. While we applied light manual filtering to remove extremely poor-quality images, some noise remains due to the inherent limitations of open-source data. Nevertheless, performance trends on CityLens-Mapillary closely mirror those on the original dataset, demonstrating that open-source platforms can still provide strong utility for socioeconomic prediction. These findings highlight the viability and promise of open-source alternatives like, especially for future research in urban socioeconomic sensing.
Table 3: Number of valid prediction tasks per indicator across regions in CityLens vs. CityLens-Mapillary.
Task Google/Baidu Mapillary
GDP 1000 805
Population 1000 824
House Price 777 527
Public Transport 500 409
Drive Ratio 500 413
Mental Health 500 416
Accessibility to Health 1000 809
Life Expectancy 193 89
Building Height 1000 813
Violent Crime 396 345
Bachelor Ratio 500 382
Table 4: Main Results on Feature-Based Regression method from Mapillary data.The values in the table represent $R^{2}$ scores. “Mean” denotes the average performance across tasks, and “SD” refers to the standard deviation. In each row, bold indicates the best result, and underline denotes the second-best.
Domain Econ. Crime Trans. Env. Health Edu. Overall
Tasks GDP Pop. HP VC PT DR BH MH AH LE BR Mean SD
Gemma3-4B 0.390 0.132 0.100 0.013 0.274 0.170 0.532 0.075 0.332 0.092 0.097 0.201 0.153
Gemma3-12B 0.453 0.148 0.146 0.020 0.348 0.242 0.573 0.039 0.342 0.174 0.061 0.231 0.171
Gemma3-27B 0.471 0.188 0.165 0.016 0.299 0.274 0.583 0.074 0.353 0.205 0.088 0.247 0.165
Qwen2.5VL-3B 0.406 0.142 0.206 0.007 0.228 0.199 0.563 0.028 0.365 0.126 0.040 0.210 0.166
Qwen2.5VL-7B 0.401 0.161 0.150 0.056 0.376 0.205 0.554 -0.189 0.357 0.043 0.160 0.207 0.196
Qwen2.5VL-32B 0.408 0.148 0.220 -0.018 0.329 0.220 0.563 -0.044 0.341 0.085 0.107 0.214 0.176
Llama4-Scout 0.401 0.161 0.053 -0.010 0.295 0.288 0.570 0.078 0.361 -0.0004 0.167 0.215 0.176
Llama4-Maverick 0.470 0.171 0.157 -0.007 0.287 0.362 0.594 0.082 0.357 -0.0004 0.088 0.233 0.188
Mistral-small-3.1-24B 0.462 0.175 0.109 0.021 0.376 0.281 0.560 0.112 0.366 -0.0004 0.099 0.233 0.179
Phi-4-multimodal 0.438 0.170 0.138 0.017 0.308 0.240 0.576 0.034 0.286 0.133 0.064 0.219 0.166
Nova-lite-v1 0.424 0.131 0.163 0.002 0.273 0.229 0.533 0.046 0.292 -0.0004 0.116 0.201 0.163
Minimax-01 0.433 0.139 0.206 -0.016 0.362 0.250 0.587 0.030 0.374 -0.0004 0.128 0.227 0.186
Gemini-2.0-Flash 0.443 0.157 0.130 -0.002 0.444 0.326 0.607 0.093 0.357 0.059 0.111 0.248 0.187
Gemini-2.5-Flash 0.452 0.210 0.198 0.040 0.362 0.246 0.598 0.004 0.297 -0.0004 0.085 0.227 0.183
GPT-4o-mini 0.394 0.135 0.151 0.007 0.273 0.371 0.533 0.071 0.384 -0.0004 0.071 0.217 0.173
GPT-4.1-mini 0.478 0.185 0.092 0.019 0.350 0.259 0.574 0.011 0.304 -0.0004 0.108 0.216 0.186
GPT-4.1-nano 0.429 0.179 0.037 0.005 0.347 0.227 0.520 0.112 0.332 0.025 0.113 0.211 0.166
A.4 Nonlinear Regressor Results under Feature-Based Regression Setting
We further conduct additional experiments using nonlinear regressors, including Random Forest, XGBoost, and MLP, to regress on the 13-attribute vectors extracted from Gemma3-27B and Llama4-Maverick. The results are summarized in the Table 5. While nonlinear regressors offer modest improvements over linear models in certain tasks such as GDP prediction, they do not consistently outperform across tasks. In particularly challenging indicators like Violent Crime and Mental Health, R² scores remain low or even negative for all nonlinear regressors, suggesting that the primary bottleneck is not the regression capacity, but rather a limitation in the expressiveness of the 13-attribute vectors extracted by current LVLMs.
Table 5: Results of nonlinear regressors, including Random Forest, XGBoost, and MLP.
Method Model GDP Pop. HP VC PT DR BH MH AH LE BR
RF Gemma3-27B 0.563 0.234 0.106 -0.335 0.477 0.587 0.618 0.181 0.302 0.240 0.256
RF Llama4-Scout 0.516 0.239 0.114 -0.063 0.443 0.448 0.575 0.207 0.304 0.176 0.266
XGBoost Gemma3-27B 0.507 0.263 0.130 -0.533 0.357 0.392 0.571 0.199 0.244 0.129 0.174
XGBoost Llama4-Scout 0.549 0.153 0.210 -0.072 0.411 0.370 0.556 0.180 0.294 0.112 0.161
MLP Gemma3-27B 0.467 0.332 0.116 0.085 0.507 0.486 0.599 -0.053 0.296 -10.632 0.218
MLP Llama4-Scout 0.413 0.247 0.192 0.005 0.317 0.402 0.610 -0.155 0.309 -1.042 0.102
A.5 Details about CityLens Dataset
A.5.1 Summary of Indicators in CityLens
@@
Under the economy domain, we cover 7 critical indicators: GDP, house price, population, median household income, poverty 100%, poverty 200% and income Gini coefficient. For GDP, we utilize a global dataset that provides GDP estimates with a spatial resolution of 1 km Wang & Sun (2022). For population, we adopt estimates from WorldPop Tatem (2017), a global demographic dataset with 1 km spatial resolution that provides consistent population counts across countries. For house price, we collect data from multiple sources tailored to each country’s context: (1) For US cities, we use the Zillow Home Value Index (ZHVI) (Zillow, 2020), available at the ZIP code level. We map these values to census tract boundaries using spatial overlays, enabling fine-grained local prediction. (2) For UK cities, we target the Middle Layer Super Output Area (MSOA) level, obtaining house price data from Han et al. (2023). (3) For Chinese cities, we collect house price data from LianJia (Platform, 2020), one of China’s largest online real estate platforms. For median household income, poverty 100%, and poverty 200%, we obtain the raw data from SafeGraph . We obtain the ground-truth values for the income Gini coefficient from Zhang et al. (2025b).
@@
In the transport domain, we include seven indicators: PMT, VMT, PTRP, VTRP, walk and bike ratio, public transport ratio, and drive ratio, following the design of Fan et al. (2023). The underlying data is sourced from Survey (2017), which provides commuting behavior statistics at the census tract level across the United States. Table 7 outlines the definitions of the seven transport related indicators.
@@
In the crime domain, we focus on two indicators: violent crime incidence and non-violent crime incidence, both defined as the number of crime occurrences per census tract. The data is collected from the official websites of individual US cities (Chicago, 2019; New York, 2019; San Francisco, 2019), which publish annual crime reports and geolocated incident-level data.
@@
For the health domain, we include 9 kinds of indicators to capture different dimensions of urban health outcomes: obesity, diabetes, cancer, no leisre-time physical activity (LPA), mental health, physical health, depression rate, accessibility to healthcare, and life expectancy. The first 7 tasks focus on the United States only, using data from “Local Data for Better Health” (PLACES, ). The Accessibility to Healthcare task is defined globally, using a dataset that quantifies walking-only travel time to the nearest healthcare facility (Weiss et al., 2020). The Life Expectancy task targets the United Kingdom, where we use data from Han et al. (2023) to obtain male life expectancy estimates at the MSOA level.
@@
In the environment domain, we consider two indicators: Carbon Emissions and Building Height. For carbon, we use global estimates from Oda & Maksyutov (2015). We use global building height data obtained from Pesaresi & Politis (2022), which provides global coverage at a spatial resolution of 100 meters.
@@
Following Liu et al. (2023b), we use the Bachelor Ratio, defined as the proportion of residents holding a bachelor’s degree or higher, as the target variable in the education domain. The ground-truth data for this indicator is obtained from SafeGraph (SafeGraph, ), which provides fine-grained demographic datasets across the United States.
@@
Economy GDP Wang & Sun (2022)
House Price Zillow (2020); Han et al. (2023)
Platform (2020)
Population Tatem (2017)
Median Income SafeGraph
Poverty 100% SafeGraph
Poverty 200% SafeGraph
Income Gini Coefficient Zhang et al. (2025b)
Education Bachelor Ratio SafeGraph
Crime Violent Chicago (2019); New York (2019) San Francisco (2019)
Non-Violent Chicago (2019); New York (2019) San Francisco (2019)
Transport PMT Survey (2017)
VMT Survey (2017)
PTRP Survey (2017)
VTRP Survey (2017)
Walk and Bike Survey (2017)
Drive Ratio Survey (2017)
Public Transport Survey (2017)
Health Obesity PLACES
Diabetes PLACES
LPA PLACES
Cancer PLACES
Mental Health PLACES
Physical Health PLACES
Depression Rate PLACES
Life Expectancy Han et al. (2023)
Accessibility to Healthcare Weiss et al. (2020)
Environment Carbon Emissions Oda & Maksyutov (2015)
Building Height Pesaresi & Politis (2022)
@@
A.5.2 Data Mapping and Aggregating
In CityLens, each region serves as a unit of prediction and is represented by 1 satellite image and 10 street view images. For each region, we associate one scalar label for the target indicator by mapping and aggregating raw tabular data from heterogeneous sources. Regarding the label mapping and aggregation strategies:
  • For Public Transport Ratio, Drive Ratio, Mental Health, and Violent Crime, we follow the methodology in Fan et al. (2023). The data is provided at the CT level, so we use CT as the key to directly map indicators to regions. For Bachelor Ratio, the original data is also available at the CT level, so we apply the same CT-based mapping.
  • For House Price:
    • In the U.S., the original values are provided by Zillow at the ZIP code level. Using the official crosswalk between ZIP codes and CTs, we assign ZIP-level values to CTs. Since some CTs and ZIP codes do not align perfectly, we use averaging in cases of overlap.
    • In China, the raw dataset consists of individual records with latitude and longitude. We aggregate values by averaging all records falling within each satellite image’s coverage.
    • In the U.K., the data is already available at the MSOA level, so we perform direct MSOA-to-region mapping.
  • For Life Expectancy, we use the same MSOA-based strategy.
  • For globally available GDP, Population, Building Height, and Accessibility to Healthcare, the original data is provided in GeoTIFF format. For each satellite image, we extract the values covering the image’s geographic extent and compute the average as the region-level indicator.
A.6 Additional Information about Three Evaluation Methodologies
A.6.1 Prompt Design and Case Analysis
We provide additional insights into the design and behavior of LVLMs under the three evaluation methodologies. Table 8 summarizes the representative prompts used in the three paradigms, highlighting their structural differences and role-setting strategies. Figure 9 shows a case in the Direct Metric Prediction setting, where the model refuses to estimate GDP due to insufficient information. Figure 10 shows an example under the Normalized Estimation setting, where the model is asked to predict the bachelor ratio based on regional images. Figure 11 presents an example from the Feature-Based Regression method, where the model scores a street view image along 13 predefined urban visual attributes to support downstream prediction.
@@ A.6.1 Prompt Design and Case Analysis
Figure 10: Case example for Normalized Metric Estimation.
Figure 11: Prompt template for guiding large vision-language models to extract 13 visual features from a street view image.
A.6.2 Supplementary Comparison of Three Evaluation Methodologies
@@
From the figures, it is evident that the Feature-Based approach, where the large vision-language model acts as a feature enhancer, consistently outperforms the Direct and Normalized approaches, in which the model is expected to behave as a numerical predictor. This suggests that current LVLMs, while powerful in perceptual and language tasks, are still more effective when used to extract structured visual representations rather than to directly generate precise socioeconomic estimates. Although large vision-language models have made impressive strides, accurately predicting fine-grained, region-level socioeconomic indicators remains highly challenging, highlighting the need for further advancements.
@@
A.7 Additional Experimental Setup Details
A.7.1 LVLMs
We consider a diverse set of LVLMs as baselines to benchmark our proposed methods. The selected models include both open-source and proprietary systems, covering a range of model sizes and capabilities. We choose Gemma3-4B/12B/27B (Team, 2025), Qwen2.5VL-3B/7B/32B (Bai et al., 2025), Llama4-Scout/Maverick (AI, 2025a), Mistral-small-3.1-24B (AI, 2025b), Phi-4-multimodal (Abouelenin et al., 2025), MiniMax-01 (Li et al., 2025), Gemini-2.0-flash/Gemini-2.5-flash (DeepMind, 2025), GPT-4o-mini (Achiam et al., 2023), GPT-4.1-mini/nano (OpenAI, 2025) and Amazon-Nova-Lite (Amazon, 2025). One thing to note is that models in the Gemini series can accept at most 10 images as input. Therefore, for this series, we use 1 satellite image and 9 street view images per region to stay within the model’s input constraints.
A.7.2 Metrics
A.7.3 Choice of Task Input
While prior work such as Fan et al. (2023) uses 20 or more street view images to represent each region, we find that this setup is often impractical for LVLMs. Specifically, we initially experiment with 20 street view images per region, but observe that this would significantly increase the computational cost, and exceed the input limits of models like Gemini, which can only process up to 10 images per inference, and also frequently hit the token length limit of other models. Therefore, we adopt a compromise of 10 images per region to ensure compatibility across models while maintaining sufficient visual context.
A.7.4 The Example of CoT Prompting
Following the designs of Zhang et al. (2025b) and Xu et al. (2024), we implement a Chain-of-Thought (CoT) prompting strategy tailored to the urban socioeconomic sensing context. Here, we present an example CoT prompt designed specifically for the House Price task.
Figure 13: CoT prompt example for the House Price task.
A.8 Alternative Output Designs and Evaluation Strategies
A.8.1 Exploring Multiple-Choice Style Answering Formats
All three evaluation formats used in CityLens can be naturally framed as regression tasks. This design choice aligns with established paradigms in recent related works Manvi et al. (2024b)Manvi et al. (2024a) and Zhang et al. (2025b), where urban socioeconomic indicators are typically formulated as continuous variables, and numerical prediction remains the primary objective.
We acknowledge that multiple-choice formats can be useful for probing certain model capabilities, such as semantic understanding or categorical reasoning. During the development of CityLens, we explore this possibility by designing a multiple-choice version of the Population prediction task. To generate negative choices, we adopt a simple yet controlled strategy: for each ground-truth population value, three distractors were sampled from a fixed pool of plausible but incorrect values—for example, 0.05, 20, and 300. Below is an example we tested using three models. Preliminary results suggest that while the multiple-choice setting reduces the output space and improves answer interpretability, it also significantly lowers task difficulty compared to free-form regression. Consequently, we ultimately opted to focus on open-ended, regression-based evaluation, which we believe more accurately reflects the complexity and realism of urban socioeconomic sensing.
This decision is supported by emerging literature in LLM evaluation. Prior studies Aidar Myrzakhan (2024) Li et al. (2024a) have identified several systematic limitations in multiple-choice-based evaluations, including selection bias, position sensitivity, and a tendency toward random guessing—especially in smaller models. These issues may lead to inflated estimates of model capability and fail to capture the inherent complexity of the task. Moreover, the multi-choice format may fail to capture the nuanced reasoning or visual understanding required for tasks like urban socioeconomic sensing. In contrast, free-form numerical prediction—despite being more challenging—offers a more direct and faithful reflection of a model’s ability to process multimodal information and generate semantically grounded outputs.
A.8.2 Street View Caption Embedding for Socioeconomic Regression
We conduct an exploratory experiment inspired by prompt design strategies from UrbanVLP Hao et al. (2025), applying them to generate captions for street view images. As illustrated in Figure 14(a), each image is accompanied by contextual information including the city name, geographic coordinates, and scores across 13 predefined visual features. This multimodal input is then fed into a large vision-language model to produce a descriptive caption of the scene.
To leverage the semantic richness of these captions, we pass the generated texts through a BERT encoder to obtain fixed-length embeddings. These embeddings are subsequently used as input features for downstream regression tasks targeting urban socioeconomic indicators. As shown in Figure 14(b), this caption-based embedding approach achieves the best performance on the population prediction task, outperforming all other methods.
((a))
Method Population $R^{2}$
Direct <0.5<-0.5
Normalized -0.1570
Feature-Based 0.2518
Caption-Embed 0.3498
((b))
Figure 14: (a) Prompt for Caption-Embed Method. (b) Population $R^{2}$ values for different methods.
A.9 Bias Audits
In CityLens, one notable issue is the underrepresentation of Global South cities in some tasks, which may pose a risk of reinforcing biases in model predictions. To clarify, for several tasks—such as Public Transport Ratio and Mental Health—Global South cities are excluded due to the unavailability of high-quality ground-truth indicators in these regions.
Building on the analysis presented in Section 3.3, we conduct a preliminary bias audit on the GDP prediction task. Specifically, we use the Global North and Global South classification provided by Wikipedia to categorize cities as follows:
  • Global North: San Francisco, New York, Tokyo, London, Paris, Sydney
  • Global South: Beijing, Shanghai, Mumbai, Moscow [^1] Note: the classification of Moscow is debated in some literature., Sao Paulo, Nairobi, Cape Town
We then compare model performance on the GDP prediction task between these two groups. As shown in the table below, we observe that models perform significantly better on Global North cities, achieving substantially higher $R^{2}$ and lower nRMSE, suggesting better predictive reliability and fit. This performance gap highlights a geographic disparity in model behavior and may point to underlying biases in how current LVLMs generalize across different socioeconomic and cultural contexts. The fact that these differences emerge under a consistent task formulation and input structure strengthens the case for further bias auditing and fairness-aware model evaluation. Additionally, in Figure 5(a), we observe that Mumbai and Moscow yield significantly negative $R^{2}$ values, further reinforcing the presence of geographic bias in model performance.
Table 9: $R^{2}$ and nRMSE for Global North and Global South Cities
City $R^{2}$ nRMSE
Global North 0.399 0.787
Global South 0.168 0.869
A.10 Hypotheses on LVLMs for Urban Socioeconomic Sensing
We propose to view visual concepts along two dimensions: low-level and high-level.
  • Low-level concepts refer to basic, concrete features such as greenery, Street Furniture, or vehicle density, which are often directly observable and recognizable by traditional machine learning models.
  • In contrast, high-level concepts—such as signs of poverty, infrastructure quality, or commercial activity—are abstract constructs that emerge from combinations of multiple visual and contextual signals. These are typically more entangled with socioeconomic meaning and are closer to human-level interpretation.
One of the key strengths of LVLMs lies in their ability to go beyond recognizing low-level features and implicitly perceive and reason about high-level visual semantics. While this ability represents a significant advancement, it also introduces major challenges—it becomes extremely difficult to analyze whether and how the model correctly interprets high-level visual features. For instance, when it comes to abstract indicators like Bachelor Ratio, it is inherently difficult to isolate a single visual factor as the dominant predictor. These outcomes are typically influenced by acombination of visual signals (e.g., signs of poverty, commercial activity, infrastructure conditions), and the relationship between them is complex and entangled.
While these models exhibit impressive capabilities in perceiving a broader spectrum of visual concepts, they also face well-known limitations, most notably hallucination and instability in perception. When identifying low-level features, LVLMs may overlook relevant signals or generate hallucinated content. At the high-level, models may fail to correctly identify or interpret key contextual cues related to the target indicator, leading to incomplete or biased reasoning. For example, we observe in the BR task that GLM-4.1v-Thinking tends to over-reason, gradually drifting away from task-relevant semantics.
The above discussion is grounded in our hypothesis about how LVLMs operate in urban socioeconomic sensing tasks: When presented with a socioeconomic prediction task, an LVLM may follow a multi-level reasoning process. It first identifies high-level visual concepts that are semantically associated with the target indicator. For example, to estimate the bachelor ratio, the model may consider abstract cues such as signs of poverty, commercial activity, and various other concepts. To support the recognition of these high-level visual concepts, the model must detect corresponding low-level visual features from the input images—such as building density, greenery coverage, and so on. These low-level features are more concrete and directly extractable from satellite or street view images. The model then aggregates these features into high-level visual concepts, which are further cross-checked and composed through reasoning to produce the final prediction.
However, if key low-level features are missed or hallucinated during recognition, the resulting high-level concept understanding may be distorted, ultimately leading to inaccurate predictions. Therefore, we believe that understanding why LVLMs struggle with urban socioeconomic sensing is a highly meaningful and non-trivial challenge that deserves deeper investigation. Moreover, the validity of our proposed hypothesis regarding the multi-level visual reasoning process in LVLMs is itself an open question.